# No Black Boxes: Interpretable and Interactable Predictive Healthcare with Knowledge-Enhanced Agentic Causal Discovery

### v1

@@ Abstract
Deep learning models trained on extensive Electronic Health Records (EHR) data have achieved high accuracy in diagnosis prediction, offering the potential to assist clinicians in decision-making and treatment planning. However, these models lack two crucial features that clinicians highly value: interpretability and interactivity. The “black-box” nature of these models makes it difficult for clinicians to understand the reasoning behind predictions, limiting their ability to make informed decisions. Additionally, the absence of interactive mechanisms prevents clinicians from incorporating their own knowledge and experience into the decision-making process. To address these limitations, we propose II-KEA, a knowledge-enhanced agent-driven causal discovery framework that integrates personalized knowledge databases and agentic LLMs. II-KEA enhances interpretability through explicit reasoning and causal analysis, while also improving interactivity by allowing clinicians to inject their knowledge and experience through customized knowledge bases and prompts. II-KEA is evaluated on both MIMIC-III and MIMIC-IV, demonstrating superior performance along with enhanced interpretability and interactivity, as evidenced by its strong results from extensive case studies.
@@
Xiaoxue Han 1, Pengfei Hu 1, Jun-En Ding 1, Chang Lu 1, Feng Liu 1, Yue Ning 1 1Stevens Institute of Technology
@@ 1 Introduction
Accurate diagnosis prediction is crucial for improving clinical outcomes by enabling timely interventions and optimizing treatment planning. In recent years, the growing availability of Electronic Health Records (EHR) (e.g., MIMIC datasets Johnson et al. (2016, 2023)) has provided valuable real-world data, allowing researchers to develop more advanced and complex deep learning models to uncover predictive patterns from a data science perspective. These models often integrate domain knowledge of medical concepts to identify intricate correlations in disease progression and comorbidities, demonstrating promising predictive performance. Despite their success in prediction accuracy, these methods have two key limitations:
  • Lack of interpretability: deep learning models inherently function as “black boxes,” offering little transparency into the clinical reasoning behind their predictions.
@@
The lack of both interpretability and interactivity undermines trust and acceptance among healthcare professionals who depend on these predictions for informed decision-making. In recent years, Large Language Models (LLMs) have demonstrated extensive knowledge, strong instruction-following capabilities, and impressive reasoning abilities, offering promising solutions to address these limitations. The development of agentic LLMs has further enhanced their flexibility and capabilities through dynamic interactions with the environment, tool utilization, and inter-agent collaboration. These advancements hold great potential for enabling clinicians to engage more effectively with predictive models, fostering greater adaptability and user-driven refinement.
@@ 2.1.1 Clinical datasets
Let $\mathcal{D}$ denote the complete set of diseases, and $\mathcal{P}_{train}$ denote the patient set from training data. The Disease Transition Probability Matrix, denoted as $\mathbf{A}_{T}\in\mathbb{R}^{\lvert\mathcal{D}\rvert\times\lvert\mathcal{D}\rvert}$, captures the probability of disease $j$ occurring after disease $i$. The underlying intuition is that temporal precedence is a necessary but not sufficient condition for causality. Identifying diseases that frequently follow a target disease helps narrow down potential causal candidates. By pre-selecting frequently co-occurring diseases, we provide the LLM agent with a shortlist of candidates, reducing its workload when assessing causal relationships, as we will discuss later. In constructing this matrix, for each visit, we define disease B as a successor of disease A if:
@@
$\mathcal{N}_{a,b}=\sum_{p\in\mathcal{P}_{\text{train}}}\sum_{i=1}^{m_{p-1}}\mathbb{I}[a\in\mathcal{D}_{p}^{i}\land(b\in\mathcal{D}_{p}^{i+1}\lor b\in\mathcal{D}_{p}^{i})],$ (2)
$\mathbb{I}[\cdot]$ is the indicator function. $\mathbf{A}_{T}[a,b]$ is an entry of $\mathbf{A}_{T}$, representing the transition probability between disease $a$ and disease $b$. $\mathcal{P}_{train}$ denotes the set of all patients in the training set, and $\mathcal{D}_{p}^{i}$ represents the set of diseases diagnosed for patient $p$ during their $i$-th visit. Note that $\mathbf{A}_{T}$ is not necessarily symmetric, meaning that $\mathbf{A}_{T}[a,b]\neq\mathbf{A}_{T}[b,a]$ in general. The diagnosis matrix $\mathbf{A}_{D}\in\mathcal{R}^{\lvert\mathcal{P}_{train}\rvert\times\lvert\mathcal{D}\rvert}$ records the occurrence of each disease for all patients. We consider the occurrence of disease $a$ for a patient to be 1 if the patient is has been diagnosed with the disease in any revisits:
@@ 2.2 Multi-agent Framework
Knowledge Synthesis Agent, $\mathcal{A}_{knowledge}$. The role of $\mathcal{A}_{knowledge}$ is to generate high-quality contextual information for the causal discovery process. Its generation process consists of two steps. In the first step, the agent is provided with the database metadata, the patient’s diagnosis history $D_{p}$, and the candidate disease set $C_{p}$. It is responsible for generating a query text to retrieve relevant information from the database. This query text should effectively summarize $D_{p}$ and $C_{p}$ while being tailored to the specific database based on its metadata, which defines its characteristics and content. We then encode the query text using the same pre-trained Sentence-BERT model and retrieve the $k$ most relevant documents. In the second step, $A_{ks}$ performs reasoning-in-documents, refining the retrieved information by removing redundancies and generating a concise summary. These summarized documents are then stored for use by the causal discovery agent, enabling a Retrieval-Augmented Generalization (RAG) approach. We summarize the workflow of the $\mathcal{A}_{knowledge}$ in Algorithm 1.
Causal Discovery Agent, $\mathcal{A}_{causal}$. The role of $\mathcal{A}_{causal}$ is to identify potential causal relationships among a set of diseases. We provide it with the patient’s diagnosis history set $D_{p}$ and the candidate disease set $C_{p}$ as a whole, along with the summarized external knowledge generated by $\mathcal{A}_{knowledge}$.
@@
    Hypothesis generation. Given the summarized external knowledge and the empty graph $\mathcal{G}^{\emptyset}$, which consists of all entities in $D_{p}$ and $C_{p}$ with no initial relations, the $\mathcal{A}_{causal}$ LLM generates an initial causal graph as a directed acyclic graph (DAG), $\mathcal{G}^{s}_{t=0}$.
@@
    Model fitting. At each iteration $t$, we fit the causal model using a data-driven approach with real-world observations. Specifically, we measure the log-likelihood, $l_{t}$, of the diagnosis matrix $\mathbf{A}_{D}$ under the current model $\mathcal{G}^{s}_{t}$.
    $\begin{split}l_{t}=\sum_{p\in\mathcal{P}_{train}}\sum_{a\in\mathcal{D}}\log P(X_{p}^{a}\mid\{X_{p}^{b}\mid b\in\mathbf{Pa}(a)\})\end{split}$ (5)
@@
Decision-Making Agent, $\mathcal{A}_{decision}$. $\mathcal{A}_{decision}$ integrates and evaluates all available information, including diagnosis history sets, summarized knowledge, and the causal graph, to make the final prediction on a patient’s diagnosis. Additionally, clinicians or users can provide their preferences, comments, or experiences to customize the prediction. For example, they may indicate that they are particularly concerned about kidney-related diseases. The agent is then tasked with producing the diagnosis list in a structured format along with an explanation of the reasoning behind its decision. We summarize the workflow of the Decision-Making Agent in Algorithm 3.
@@ 2.3 II-KEA Inference
II-KEA does not involve any training process but requires data preprocessing. First, the EHR training dataset is processed to construct the Disease Transition Probability Matrix $\mathbf{A}_{T}$ and the Diagnosis Matrix $\mathbf{A}_{D}$, as described in Section 2.1.1. Additionally, a knowledge vector database $\Gamma$ is prepared following Section 2.1.2. Both matrices and the database are stored for later inference. During inference, for each patient, we collect their diagnosis history $\mathcal{D}_{p}$ and apply the preprocessing steps outlined in Section 2.1.1 to determine the candidate disease set $\mathcal{S}_{p}$. The Knowledge Synthesis Agent $\mathcal{A}_{knowledge}$ then retrieves relevant documents from $\Gamma$ and summarizes them into $\Gamma_{p}^{summary}$. Next, the Causal Discovery Agent $\mathcal{A}_{causal}$ iteratively uncovers causal relationships within the expanded disease set $\mathcal{D}_{p}\cup\mathcal{S}_{p}$, leveraging both external knowledge from $\Gamma_{p}^{summary}$ and observational data from $\mathbf{A}_{D}$. This process results in a causal graph $\mathcal{G}^{s}$. Finally, the Decision-Making Agent $\mathcal{A}_{decision}$ integrates all available information, including the diagnosis history $\mathcal{D}_{p}$, candidate diseases $\mathcal{S}_{p}$, summarized documents $\Gamma_{p}^{summary}$, causal graph $\mathcal{G}^{s}$, and an optional clinician-provided comment $\mathcal{C}$. Using this information, the model predicts the patient’s diagnosis and provides explanations for the decision. We provide the overview of II-KEA in Figure 2, the prompt details in Appendix D. The workflow of its inference process is shown in Algorithm 4.
@@ 3.3 Baselines
Table 1: Prediction Results on MIMIC-III and MIMIC-IV for Diagnosis Prediction. The best results for each metric are highlighted, and the second bests results are underlined.
@@ 4 Related Work
We categorize prior work into clinical prediction (section 4.1), agentic approaches (section 4.2), and causal inference (appendix B.1).
@@ 4.1 Predictive Healthcare in EHR
Predictive modeling in healthcare has advanced significantly with the adoption of deep learning techniques applied to Electronic Health Records (EHR) data. Existing neural network-based models, including RNN/Attention-based approaches (Choi et al., 2016a, 2017; Ma et al., 2020), graph-based models (Choi et al., 2017; Ma et al., 2018; Lu et al., 2021), and Transformer-based architectures (Shang et al., 2019b; Luo et al., 2020; Poulain and Beheshti, 2024), have demonstrated effectiveness in capturing temporal patterns and interactions among medical concepts. Recent work (Jiang et al., 2024) has explored leveraging external knowledge sources beyond hierarchical structures such as ICD-9-CM by integrating Large Language Models (LLMs) to enhance medical predictions.
Still, most models remain black boxes, offering limited interpretability and restricting healthcare professionals from interacting with the system to refine or adjust predictions. In clinical applications, predictive models must provide faithful explanations, such as causal pathways, and allow interactive refinement based on expert guidance.
@@ 5 Conclusion
In this paper, we introduce II-KEA, a knowledge-enhanced Agentic Causal Discovery framework designed for interpretable and interactive diagnosis prediction. II-KEA consists of three LLM-based agents working collaboratively and is powered by both clinical datasets and domain knowledge. We evaluate II-KEA on the MIMIC-III and MIMIC-IV datasets and conduct both ablation and case studies to demonstrate its effectiveness. The ethical consideration can be checked in Appendix A.
@@ 6 Limitation and Future Work
II-KEA showcases a promising paradigm for interpretable and interactive diagnosis prediction by leveraging LLM agents. However, II-KEA has significant potential to solve a broader range of medical challenges. Future work includes:
@@
A Privacy and Ethical Statement
B Additional Related Work
B.1 Causality Inference on LLM
C Pseudocodes
Algorithm 1 Knowledge Synthesis Agent, $\mathcal{A}_{knowledge}$.

### v2

@@ Abstract
Deep learning models trained on extensive Electronic Health Records (EHR) data have achieved high accuracy in diagnosis prediction, offering the potential to assist clinicians in decision-making and treatment planning. However, these models lack two crucial features that clinicians highly value: interpretability and interactivity. The “black-box” nature of these models makes it difficult for clinicians to understand the reasoning behind predictions, limiting their ability to make informed decisions. Additionally, the absence of interactive mechanisms prevents clinicians from incorporating their own knowledge and experience into the decision-making process. To address these limitations, we propose II-KEA, a knowledge-enhanced agent-driven causal discovery framework that integrates personalized knowledge databases and agentic LLMs. II-KEA enhances interpretability through explicit reasoning and causal analysis, while also improving interactivity by allowing clinicians to inject their knowledge and experience through customized knowledge bases and prompts. II-KEA is evaluated on both MIMIC-III and MIMIC-IV, demonstrating superior performance along with enhanced interpretability and interactivity, as evidenced by its strong results from extensive case studies. Our code is available at https://github.com/hanxiaoxue114/II-KEA_HealthcareAgent.
@@ 1 Introduction
Accurate diagnosis prediction is crucial for improving clinical outcomes by enabling timely interventions and optimizing treatment planning. In recent years, the growing availability of Electronic Health Records (EHR) (e.g., MIMIC datasets Johnson et al. (2016, 2023)) has provided valuable real-world data, allowing researchers to develop more advanced and complex deep learning models Lu et al. (2021, 2022); Poulain and Beheshti (2024); Jiang et al. (2024); Hu et al. (2024) to uncover predictive patterns from a data science perspective. These models often integrate domain knowledge of medical concepts to identify intricate correlations in disease progression and comorbidities, demonstrating promising predictive performance. Despite their success in prediction accuracy, these methods have two key limitations:
  • Lack of interpretability: deep learning models inherently function as “black boxes”, offering little transparency into the clinical reasoning behind their predictions.
@@
The lack of both interpretability and interactivity undermines trust and acceptance among healthcare professionals who depend on these predictions for informed decision-making. In recent years, Large Language Models (LLMs) have demonstrated extensive knowledge, strong instruction-following capabilities, and impressive reasoning abilities, offering promising solutions to address these limitations. The development of agentic LLMs Wang et al. (2024); Kim et al. (2024); Zuo et al. (2024) has further enhanced their flexibility and capabilities through dynamic interactions with the environment, tool utilization, and inter-agent collaboration. These advancements hold great potential for enabling clinicians to engage more effectively with predictive models, fostering greater adaptability and user-driven refinement.
@@ 2.1.1 Clinical datasets
Let $\mathcal{D}$ denote the complete set of diseases, and $\mathcal{P}_{tr}$ denote the patient set from training data. The Disease Transition Probability Matrix, denoted as $\mathbf{A}_{T}\in\mathbb{R}^{\lvert\mathcal{D}\rvert\times\lvert\mathcal{D}\rvert}$, captures the probability of disease $j$ occurring after disease $i$. The underlying intuition is that temporal precedence is a necessary but not sufficient condition for causality. Identifying diseases that frequently follow a target disease helps narrow down potential causal candidates. By pre-selecting frequently co-occurring diseases, we provide the LLM agent with a shortlist of candidates, reducing its workload when assessing causal relationships, as we will discuss later. In constructing this matrix, for each visit, we define disease B as a successor of disease A if:
@@
$\mathcal{N}_{a,b}=\sum_{p\in\mathcal{P}_{\text{tr}}}\sum_{i=1}^{m_{p-1}}\mathbb{I}[a\in\mathcal{D}_{p}^{i}\land(b\in\mathcal{D}_{p}^{i+1}\lor b\in\mathcal{D}_{p}^{i})],$ (2)
and $\mathbb{I}[\cdot]$ is the indicator function. $\mathbf{A}_{T}[a,b]$ is an entry of $\mathbf{A}_{T}$, representing the transition probability between disease $a$ and disease $b$. $\mathcal{P}_{tr}$ denotes the set of all patients in the training set, and $\mathcal{D}_{p}^{i}$ represents the set of diseases diagnosed for patient $p$ during their $i$-th visit. Note that $\mathbf{A}_{T}$ is not necessarily symmetric, meaning that $\mathbf{A}_{T}[a,b]\neq\mathbf{A}_{T}[b,a]$ in general. The diagnosis matrix $\mathbf{A}_{D}\in\mathcal{R}^{\lvert\mathcal{P}_{tr}\rvert\times\lvert\mathcal{D}\rvert}$ records the occurrence of each disease for all patients. We consider the occurrence of disease $a$ for a patient to be 1 if the patient is has been diagnosed with the disease in any revisits:
@@ 2.2 Multi-agent Framework
Knowledge Synthesis Agent, $\mathcal{A}_{\text{knowledge}}$. The role of $\mathcal{A}_{\text{knowledge}}$ is to generate high-quality contextual information for the causal discovery process. Its generation process consists of two steps. In the first step, the agent is provided with the database metadata, the patient’s diagnosis history $D_{p}$, and the candidate disease set $C_{p}$. It is responsible for generating a query text to retrieve relevant information from the database. This query text should effectively summarize $D_{p}$ and $C_{p}$ while being tailored to the specific database based on its metadata, which defines its characteristics and content. We then encode the query text using the same pre-trained Sentence-BERT model and retrieve the $k$ most relevant documents. In the second step, $A_{ks}$ performs reasoning-in-documents, refining the retrieved information by removing redundancies and generating a concise summary. These summarized documents are then stored for use by the causal discovery agent, enabling a Retrieval-Augmented Generalization (RAG) approach. We summarize the workflow of the $\mathcal{A}_{\text{knowledge}}$ in Algorithm 1.
Causal Discovery Agent, $\mathcal{A}_{\text{causal}}$. The role of $\mathcal{A}_{\text{causal}}$ is to identify potential causal relationships among a set of diseases. We provide it with the patient’s diagnosis history set $D_{p}$ and the candidate disease set $C_{p}$ as a whole, along with the summarized external knowledge generated by $\mathcal{A}_{\text{knowledge}}$.
@@
    Hypothesis generation. Given the summarized external knowledge and the empty graph $\mathcal{G}^{\emptyset}$, which consists of all entities in $D_{p}$ and $C_{p}$ with no initial relations, the $\mathcal{A}_{\text{causal}}$ LLM generates an initial causal graph as a directed acyclic graph (DAG), $\mathcal{G}^{s}_{t=0}$.
@@
    Model fitting. At each iteration $t$, we fit the causal model using a data-driven approach with real-world observations. Specifically, we measure the fitting score as the log-likelihood, $l_{t}$, of the diagnosis matrix $\mathbf{A}_{D}$ under the current model $\mathcal{G}^{s}_{t}$.
    $\begin{split}l_{t}=\sum_{p\in\mathcal{P}_{tr}}\sum_{a\in\mathcal{D}}\log P(X_{p}^{a}\mid\{X_{p}^{b}\mid b\in\mathbf{Pa}(a)\})\end{split}$ (5)
@@
Decision-Making Agent, $\mathcal{A}_{\text{dm}}$. The role of $\mathcal{A}_{\text{dm}}$ is to integrate and evaluate all available information, including diagnosis history sets, summarized knowledge, and the causal graph, to make the final prediction on a patient’s diagnosis. Additionally, clinicians or users can provide their preferences, comments, or experiences to customize the prediction. For example, they may indicate that they are particularly concerned about kidney-related diseases. The agent is then tasked with producing the diagnosis list in a structured format along with an explanation of the reasoning behind its decision. We summarize the workflow of the Decision-Making Agent in Algorithm 3.
@@ 2.3 II-KEA Inference
II-KEA does not involve any training process but requires data preprocessing. First, the EHR training dataset is processed to construct the Disease Transition Probability Matrix $\mathbf{A}_{T}$ and the Diagnosis Matrix $\mathbf{A}_{D}$, as described in Section 2.1.1. Additionally, a knowledge vector database $\Gamma$ is prepared following Section 2.1.2. Both matrices and the database are stored for later inference. During inference, for each patient, we collect their diagnosis history $\mathcal{D}_{p}$ and apply the preprocessing steps outlined in Section 2.1.1 to determine the candidate disease set $\mathcal{S}_{p}$. The Knowledge Synthesis Agent $\mathcal{A}_{\text{knowledge}}$ then retrieves relevant documents from $\Gamma$ and summarizes them into $\Gamma_{p}^{\text{summary}}$. Next, the Causal Discovery Agent $\mathcal{A}_{\text{causal}}$ iteratively uncovers causal relationships within the expanded disease set $\mathcal{D}_{p}\cup\mathcal{S}_{p}$, leveraging both external knowledge from $\Gamma_{p}^{summary}$ and observational data from $\mathbf{A}_{D}$. This process results in a causal graph $\mathcal{G}^{s}$. Finally, the Decision-Making Agent $\mathcal{A}_{\text{dm}}$ integrates all available information, including the diagnosis history $\mathcal{D}_{p}$, candidate diseases $\mathcal{S}_{p}$, summarized documents $\Gamma_{p}^{summary}$, causal graph $\mathcal{G}^{s}$, and an optional clinician-provided comment $\mathcal{C}$. Using this information, the model predicts the patient’s diagnosis and provides explanations for the decision. We provide the overview of II-KEA in Figure 2, the prompt details in Appendix C. The workflow of its inference process is shown in Algorithm 4.
@@ 3.3 Baselines
Table 1: Prediction Results on MIMIC-III and MIMIC-IV for Diagnosis Prediction. The best results for each metric are bold, and the second bests results are highlighted in gray.
@@ 4 Related Work
We categorize prior work into clinical prediction (section 4.1), agentic approaches (section 4.2), and causal inference (section 4.3).
@@ 4.1 Predictive Healthcare in EHR
Predictive modeling in healthcare has advanced significantly with the adoption of deep learning techniques (Badawy et al., 2023; Zhang et al., 2023) applied to EHR data. Existing neural network-based models, including RNN/Attention-based approaches (Choi et al., 2016a, 2017; Ma et al., 2020), graph-based models (Choi et al., 2017; Ma et al., 2018; Lu et al., 2021), and Transformer-based architectures (Shang et al., 2019b; Luo et al., 2020; Poulain and Beheshti, 2024), have demonstrated effectiveness in capturing temporal patterns and interactions among medical concepts. Recent work (Jiang et al., 2024) has explored leveraging external knowledge sources beyond hierarchical structures such as ICD-9-CM by integrating Large Language Models (LLMs) to enhance medical predictions.
Still, most models remain black boxes (Zhang et al., 2024), offering limited interpretability and restricting healthcare professionals from interacting with the system to refine or adjust predictions. In clinical applications, predictive models must provide faithful explanations, such as causal pathways, and allow interactive refinement based on expert guidance.
@@ 4.2 LLM Agents for Healthcare AI
4.3 Causality Inference on LLM
5 Computation Cost
We evaluated the cost of performing one prediction loop with II-KEA using a sample of 10 patients. On average, each patient required 7,803 tokens (4,871 input tokens and 2,932 output tokens), resulting in an estimated cost of approximately 0.00263 USD per patient when using GPT-4o Mini (at rates of 0.0003 USD per 1,000 input tokens and 0.0004 USD per 1,000 output tokens). The average processing time was 87 seconds per patient.
6 Privacy and Ethical Statement
@@ 7 Conclusion
In this paper, we introduce II-KEA, a knowledge-enhanced Agentic Causal Discovery framework designed for interpretable and interactive diagnosis prediction. II-KEA consists of three LLM-based agents working collaboratively and is powered by both clinical datasets and domain knowledge. We evaluate II-KEA on the MIMIC-III and MIMIC-IV datasets and conduct both ablation and case studies to demonstrate its effectiveness. The ethical consideration can be checked in Appendix 6.
@@ 8 Limitation and Future Work
II-KEA showcases a promising paradigm for interpretable and interactive diagnosis prediction by leveraging LLM agents. While II-KEA focuses on diagnosis prediction, future research will explore to solve a broader range of medical challenges tailored to clinicians’ needs, such as treatment planning and personalized medical recommendations. Future work includes:
@@
Acknowledgment
We thank the anonymous reviewers and area chairs for their valuable feedback. We used ChatGPT to polish our writing. This work was supported in part by the U.S. National Science Foundation under grants 2047843 and 2437621. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation.
@@ Acknowledgment
A Potential Risk of AI clinical prediction
There are three kinds of potential risks by using agentic AI for predictive healthcare:
(1) Clinical Safety Risks. AI clinical prediction systems pose significant clinical safety risks including diagnostic errors from overreliance on black-box models, perpetuation of healthcare biases present in training data, and poor generalization to populations underrepresented in datasets. These issues can lead to differential prediction accuracy across demographic groups and unreliable performance in real-world clinical scenarios.
(2) Ethical & Legal Risks. Ethical and legal concerns arise from the lack of interpretability in traditional AI systems, creating accountability challenges when predictions contribute to adverse outcomes. Additionally, opaque AI decision-making may compromise informed consent and patient autonomy, while large-scale EHR data requirements raise privacy and security concerns regarding unauthorized access to sensitive medical information.
(3) Operational Risks. Operational challenges include disruption of established clinical workflows, potential skill atrophy among healthcare professionals who become overly dependent on AI recommendations, and false confidence generated by systems that provide predictions without uncertainty quantification. These issues particularly affect trainees and may lead to inadequate consideration of alternative diagnoses.
The II-KEA framework mitigates these risks through explicit causal reasoning and detailed explanations that enable clinician evaluation of AI recommendations, reducing overreliance and supporting accountability. Its interactive design requires active clinician participation, preserving clinical judgment while allowing healthcare professionals to inject domain knowledge and maintain diagnostic control. The multi-agent architecture provides validation layers, external knowledge integration reduces bias, and transparent reasoning supports both audit trails and informed consent processes.
B Pseudocodes
The pseudocodes of three agents in II-KEA are demonstrated in Algorithms 1 to 3, and the Algorithm 4 shows the overall process of II-KEA.
Algorithm 1 Knowledge Synthesis Agent, $\mathcal{A}_{\text{knowledge}}$.